Top Company holds 90 percent of Bottom Company’s common stock. In the current year, Top reports sales of $830,000 and cost of goods sold of $622,500. For this same period, Bottom has sales of $357,000 and cost of goods sold of $249,900.

During the current year, Bottom sold merchandise to Top for $188,000.

The parent still possesses 35 percent of this inventory at the current year-end. Bottom had established the transfer price based on its normal gross profit rate. What are the consolidated sales and cost of goods sold?
Answer #1

Ending unrealized gross profit = 188000*35% *(830000-622500)/830000 = $16,450

Consolidated cost of goods sold = 622500+24900-188000+16450 = $475,850

Consolidated sale = 830000+357000-188000 = $999,000